# Confidence on the Focal: Conformal Prediction with Selection-Conditional Coverage

### v1

@@ 4.1.3 Selection based on calibration quantiles
Similar to the previous two cases, directly computing ^$\widehat{{\mathcal{R}}}_{\textnormal{cq}}$ reduces the computation complexity of the reference set to $O(\max(m,n))$. As a side note, Bao et al., (2024) also considered the selection rule based on calibration quantiles (as an instance of “calibration-assisted selection”), for which they provide asymptoically FCR control guarantees under additional assumptions.
4.2 Selection based on conformal p-values
Suppose we wish to select the test units whose outcomes satisfy certain conditions. For example, Jin and Candès, 2023b studies the problem of selecting test units with large outcomes, where they formulate the task as hypothesis testing problems. Following their formulation, we consider for each $j\in[m]$ the following null and alternative hypotheses:
$H_{j,0}:Y_{n+j}\leq c_{n+j}\text{ v.s. }H_{j,1}:Y_{n+j}>c_{n+j},$ (16)
where $c_{n+j}$ is a (potentially random) threshold. When $H_{j,0}$ is rejected, we select the $j$-th test unit. For testing $H_{j,0}$in and Candès, 2023b devise the conformal p-value $p_{j}$ such that
$\mathbb{P}(p_{j}\leq t,Y_{n+j}\leq c_{j})\leq t,\text{ for any }t\in[0,1].$
In this paper, we follow their strategy, but work with a slightly different form of the conformal p-value that is uniformly more powerful than the original version in Jin and Candès, 2023b . Throughout Section 4.2, we assume the score function $S(x,y)$ to be monotonically nonincreasing in $y$, e.g., $s(x,y)=\widehat{\mu}(x)-y$. For each $i\in{\mathcal{I}}_{\textnormal{calib}}\cup{\mathcal{I}}_{\textnormal{test}}$, define $\widehat{S}_{i}=S(X_{i},c_{i})$, and for any $j\in[m]$, we define the conformal p-value as
$p_{j}=\frac{1+\sum_{i\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{\widehat{S}_{i}\geq\widehat{S}_{n+j},Y_{i}\leq c_{i}\}}{n+1}.$ (17)
We provide a proof for the validity of $p_{j}$ for testing $H_{j,0}$, as well as its comparison with Jin and Candès, 2023b in Appendix C.1 for completeness.
With these p-values, we can select the units whose conformal p-values are smaller than a threshold. In terms of the threshold choice, we consider two scenarios:
  1. 1.
    Fixed threshold: select test units whose conformal p-values (17) are below a fixed threshold $q\in(0,1)$, such as Bonferroni correction.
  2. 2.
    Data-driven threshold: select test units whose conformal p-values (17) are below a data-dependent threshold given by the Benjamini-Hochberg (BH) procedure (Benjamini and Hochberg,, 1995).
4.2.1 Fixed threshold
We first consider $\widehat{{\mathcal{S}}}_{\textnormal{fixed}}=\{j\in[m]:p_{j}\leq q\}$ with a threshold $q\in(0,1)$. We will see that there are only two potential forms of the reference set ^$\widehat{{\mathcal{R}}}_{n+j}(y)$ when varying $y\in\mathbb{R}$. For $k\in\{0,1\}$, define
^$\widehat{{\mathcal{R}}}_{n+j}^{\textnormal{fixed},k}=\leavevmode\nobreak$$\Big{\{}i\in{\mathcal{I}}_{\textnormal{calib}}:Y_{i}\leq c_{i},\widehat{S}_{i}\geq T_{j}^{\textnormal{fixed}}(k,0),\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}\geq T_{j}^{\textnormal{fixed}}(k,0)\}\in\mathfrak{S}\Big{\}}$
$\cup\Big{\{}i\in{\mathcal{I}}_{\textnormal{calib}}:Y_{i}>c_{i},\widehat{S}_{i}\geq T_{j}^{\textnormal{fixed}}(k,1),\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}\geq T_{j}^{\textnormal{fixed}}(k,1)\}\in\mathfrak{S}\Big{\}},$ (18)
where we recall that $\widehat{S}_{n+\ell}^{{\textnormal{swap}(i,j)}}$ is the score in the ($n+\ell$)-th position after swapping $(i,j)$. For $k,\ell\in\{0,1\}$,
$T^{\textnormal{fixed}}_{j}(k,\ell)=\inf\bigg{\{}t\in{\mathcal{T}}:\ell+\sum_{i\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{\widehat{S}_{i}\geq t,Y_{i}\leq c_{i}\}+k\cdot\mathds{1}\{\widehat{S}_{n+j}\geq t\}\leq q(n+1)\bigg{\}}.$ (19)
Proposition 6 presents the simplified expression for ^$\widehat{\mathcal{R}}_{n+j}(y)$ for any selected unit and any selection taxonomy, whose proof is deferred to Appendix B.7.
Proposition 6.
Suppose the selection set contains the units whose conformal p-values (17) are below a fixed value $q\in(0,1)$, i.e., $\widehat{{\mathcal{S}}}=\widehat{{\mathcal{S}}}_{\textnormal{fixed}}=\{j\in[m]\colon p_{j}\leq q\}$. Then for any selection taxonomy $\mathfrak{S}$ and any $j\in[m]$ such that $j\in\widehat{{\mathcal{S}}}_{\textnormal{fixed}}$ and $\widehat{{\mathcal{S}}}_{\textnormal{fixed}}\in\mathfrak{S}$, it holds that ^$\widehat{{\mathcal{R}}}_{n+j}(y)=\widehat{{\mathcal{R}}}^{\textnormal{fixed},0}_{n+j}$ when $y>c_{n+j}$, and ^$\widehat{{\mathcal{R}}}_{n+j}(y)=\widehat{{\mathcal{R}}}^{\textnormal{fixed},1}_{n+j}$ when $y\leq c_{n+j}$, for any $y\in\mathbb{R}$.
Finally, the JOMI prediction interval is therefore
$\widehat{{\mathcal{C}}}_{\alpha}(X_{n+j})=\big{\{}y\in{\mathcal{Y}}:y>c_{n+j},V(X_{n+j},y)\leq\widehat{q}_{0}\big{\}}\cup\big{\{}y\in{\mathcal{Y}}:y\leq c_{n+j},V(X_{n+j},y)\leq\widehat{q}_{1}\big{\}},$
where $\widehat{q}_{k}={\textnormal{Quantile}}(\{V_{i}\}_{i\in\widehat{{\mathcal{R}}}^{\textnormal{fixed},k}_{n+j}}\cup\{\infty\};1-\alpha)$ for $k=0,1$. See Algorithm LABEL:alg:pval_fix for a summary.
The computation complexity of Algorithm LABEL:alg:pval_fix is polynomial in $n$ and $m$. Note that for each $j$, computing $T_{j}^{\textnormal{fixed}}(k,\ell)$ costs $O(m)$, computing ^$\widehat{\mathcal{R}}_{n+j}^{\textnormal{fixed},k}$ costs $O(mn)$, and computing $\widehat{C}_{\alpha,n+j}$ costs $O(n)$.
Input: Calibration data ${\mathcal{D}}_{\textnormal{calib}}$; test data ${\mathcal{D}}_{{\textnormal{test}}}$; miscoverage level $\alpha$; selection taxonomy $\mathfrak{S}$;
selection rule ${\mathcal{S}}$; selection score function $S(\cdot,\cdot)$; nonconformity score function $V(\cdot,\cdot)$.
for $j\in\widehat{\mathcal{S}}={\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})$ do
       Compute $T_{j}^{\textnormal{fixed}}(k,\ell)$ as in (19) for $(k,\ell)\in\{0,1\}$. Compute ^$\widehat{{\mathcal{R}}}_{n+j}^{\text{fixed},0}$ and ^$\widehat{{\mathcal{R}}}_{n+j}^{\text{fixed},1}$ as (4.2.1). Compute $\widehat{q}_{k}=\text{Quantile}(1-\alpha;\{V_{i}\}_{i\in\widehat{\mathcal{R}}_{n+j}^{\text{fixed},k}}\cup\{\infty\})$, $k=0,1$. Compute $\widehat{C}_{\alpha,n+j}=\{y\in{\mathcal{Y}}\colon y>c_{n+j},V(X_{n+j},y)\leq\widehat{q}_{0}\}\cup\{y\in{\mathcal{Y}}\colon y\leq c_{n+j},V(X_{n+j},y)\leq\widehat{q}_{1}\}$
end for
Output: $\{\widehat{C}_{\alpha,n+j}\}_{j\in\widehat{{\mathcal{S}}}}$.
Algorithm 2 JOMI for conformal p-values with fixed threshold
4.2.2 Data-driven p-value threshold
According to Jin and Candès, 2023b , applying the BH procedure (Benjamini and Hochberg,, 1995) to the p-values $\{p_{j}\}_{j=1}^{m}$ yields a selection set with controlled false discovery rate (FDR). Formally, let $p_{(1)}\leq p_{(2)}\leq\ldots\leq p_{(m)}$ be the order statistics of the conformal p-values. The BH selection set $\widehat{{\mathcal{S}}}_{\textnormal{BH}}$ is
$\widehat{{\mathcal{S}}}_{\textnormal{BH}}=\{j\colon p_{j}\leq k^{*}q/m\},\quad k^{*}=\max\Big{\{}k:p_{(k)}\leq kq/m\Big{\}},$
with the convention that $\max\varnothing=0$.
For the BH selection rule, the reference set can also be simplified to two cases. For $k=0,1$, define
^$\widehat{{\mathcal{R}}}_{n+j}^{\textnormal{BH},k}=\leavevmode\nobreak$$\Big{\{}i\in{\mathcal{I}}_{\textnormal{calib}}:Y_{i}\leq c_{i},\widehat{S}_{i}\geq T_{j}^{\textnormal{BH}}(k,0),\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}\geq T_{j}^{\textnormal{BH}}(k,0)\}\in\mathfrak{S}\Big{\}}$
$\cup\Big{\{}i\in{\mathcal{I}}_{\textnormal{calib}}:Y_{i}>c_{i},\widehat{S}_{i}\geq T_{j}^{\textnormal{BH}}(k,1),\{\ell\in[m]:\widehat{S}_{n+\ell}^{\textnormal{swap}(i,j)}\geq T_{j}^{\textnormal{BH}}(k,1)\}\in\mathfrak{S}\Big{\}},$ (20)
and for $k,\ell=0,1$, we define
$T^{\textnormal{BH}}_{j}(k,\ell)=\inf\bigg{\{}t\in{\mathcal{T}}:\frac{\ell+\sum_{i\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{\widehat{S}_{i}\geq t,Y_{i}\leq c_{i}\}+k\cdot\mathds{1}\{\widehat{S}_{n+j}\geq t\}}{1+\sum_{j\in{\mathcal{I}}_{j}^{c}}\mathds{1}\{\widehat{S}_{\ell}\geq T_{j}^{(k,\ell)}\}}\cdot\frac{m}{n+1}\leq q\bigg{\}}.$ (21)
We prove the equivalence in the next proposition, deferring the proof to Appendix B.8.
Proposition 7.
For the BH selection rule, for any $j\in[m]$ and any selection taxonomy $\mathfrak{S}$ such that $j\in\widehat{{\mathcal{S}}}$ and $\widehat{{\mathcal{S}}}\in\mathfrak{S}$, the reference set obeys ^$\widehat{{\mathcal{R}}}_{n+j}(y)={\mathcal{R}}_{n+j}^{\textnormal{BH},1}$ when $y\leq c_{n+j}$, and ^$\widehat{{\mathcal{R}}}_{n+j}(y)={\mathcal{R}}_{n+j}^{\textnormal{BH},0}$ when $y>c_{n+j}$ as defined in (4.2.2) for any $y\in\mathbb{R}$.
Therefore, the JOMI prediction set is of a similar form as the fixed-threshold case. We summarize the entire workflow in Algorithm 3. Similarly, the computation cost of $\widehat{C}_{\alpha,n+j}$ is $O(mn)$.
Input: Calibration data ${\mathcal{D}}_{\textnormal{calib}}$; test data ${\mathcal{D}}_{{\textnormal{test}}}$; miscoverage level $\alpha$; selection taxonomy $\mathfrak{S}$;
selection rule ${\mathcal{S}}$; selection score function $S(\cdot,\cdot)$; nonconformity score function $V(\cdot,\cdot)$.
for $j\in\widehat{\mathcal{S}}={\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})$ do
       Compute $T_{j}^{\textnormal{BH}}(k,\ell)$ as in (21) for $(k,\ell)\in\{0,1\}$. Compute ^$\widehat{{\mathcal{R}}}_{n+j}^{\textnormal{BH},0}$ and ^$\widehat{{\mathcal{R}}}_{n+j}^{\textnormal{BH},1}$ as (4.2.2). Compute $\widehat{q}_{k}=\text{Quantile}(1-\alpha;\{V_{i}\}_{i\in\widehat{\mathcal{R}}_{n+j}^{\textnormal{BH},k}}\cup\{\infty\})$, $k=0,1$. Compute $\widehat{C}_{\alpha,n+j}=\{y\in{\mathcal{Y}}\colon y>c_{n+j},V(X_{n+j},y)\leq\widehat{q}_{0}\}\cup\{y\in{\mathcal{Y}}\colon y\leq c_{n+j},V(X_{n+j},y)\leq\widehat{q}_{1}\}$
end for
Output: $\{\widehat{C}_{\alpha,n+j}\}_{j\in\widehat{{\mathcal{S}}}}$.
Algorithm 3 JOMI for conformal p-values with BH-threshold
4.3 Selection based on conformal prediction sets
Practitioners may also select test unit(s) based on the property of a preliminary prediction set. For example, they may select units whose prediction intervals are shorter than a threshold (Sokol et al.,, 2024), or whose prediction sets that entirely lie above a threshold. Note that the original prediction intervals are no longer valid conditional on being selected (Jin and Candès, 2023b, ). In this section, we instantiate our general framework to re-calibrate prediction sets for the units selected in such a way.
We thus consider two stages of prediction set construction: the one constructed in the first stage, which we call the preliminary prediction set, is used for determining the selection set $\widehat{\mathcal{S}}$, while that in the second stage, which we call the selective prediction set, is the one we are to build with JOMI.
Recall the SCP framework in Section 3.1. Let $S(x,y)$ and $V(x,y)$ be the nonconformity score for the two stages, respectively. The level $1-\beta$ preliminary prediction set for the $j$-th test unit is
$\widehat{C}^{\text{prelim}}_{\beta}(X_{n+j})=\big{\{}y\in{\mathcal{Y}}:S(X_{n+j},y)\leq\eta\big{\}},$ (22)
where $\eta$ is the $K\,:=\,\lceil(1-\beta)(n+1)\rceil$-th smallest element in $\{S_{i}\}_{i=1}^{n}$.
By (22), the preliminary prediction set is fully determined by $X_{n+j}$ and $\eta$. We can thus consider a general selection rule ${\mathcal{L}}:{\mathcal{X}}\times\mathbb{R}\mapsto\{0,1\}$, where ${\mathcal{L}}(X_{n+j},\eta)=1$ means selecting the unit. For instance, selecting prediction intervals shorter than a threshold $\lambda$ gives ${\mathcal{L}}(x,\eta)=\mathds{1}\{\widehat{q}_{U}(x)-\widehat{q}_{L}(x)+2\eta\leq\lambda\}$ if we use CQR for $\widehat{C}_{\beta}^{\text{prelim}}$ (Romano et al.,, 2019), i.e., setting $S(x,y)=\max\{\widehat{q}_{L}(x)-y,y-\widehat{q}_{U}(x)\}$, where $\widehat{q}_{L}(x)$ and $\widehat{q}_{U}(x)$ are estimates of some lower and upper conditional quantiles. As another example, for binary $Y$, we might want to select the units whose prediction set is a singleton, then the selection rule can be written as ${\mathcal{L}}(x,\eta)=\mathds{1}\big{\{}S(x,1)\leq\eta<S(x,0)\text{ or }S(x,0)\leq\eta<S(x,1)\big{\}}$.
The selection set for a selection rule ${\mathcal{L}}$ is
$\widehat{{\mathcal{S}}}_{\textnormal{ps}}=\big{\{}j\in[m]:{\mathcal{L}}(X_{n+j},\eta)=1\big{\}}.$
We are to derive a general form for a slightly conservative version of the JOMI prediction set, which nevertheless has tight coverage in all our numerical experiments (see Section 6 for examples). We only present the results for $\mathfrak{S}=2^{[m]}$ for the ease of exposition. Formally, we define
$\widehat{C}^{\textnormal{ps}}_{\alpha,n+j}:=\leavevmode\nobreak$$\{y:\eta^{-}\leq S(X_{n+1},y)\leq\eta^{+}\}$ (23)
$\cup\{y:V(X_{n+j},y)\leq q_{1},S(X_{n+1},y)<\eta^{-})\}$
$\cup\{y:V(X_{n+j},y)\leq q_{2},S(X_{n+1},y)>\eta^{+})\}$
$q_{1}:={\textnormal{Quantile}}\Big{(}\big{\{}V_{i}\colon i\in[n],S_{i}\leq\eta^{-},{\mathcal{L}}(X_{i},\eta)=1\big{\}}\cup\big{\{}V_{i}:S_{i}>\eta^{-},\leavevmode\nobreak\ {\mathcal{L}}(X_{i},\eta^{-})=1\big{\}}\Big{)};$
$q_{2}:={\textnormal{Quantile}}\Big{(}\big{\{}V_{i}\colon i\in[n],S_{i}\leq\eta,{\mathcal{L}}(X_{i}.\eta^{+})=1\big{\}}\cup\big{\{}V_{i}:S_{i}>\eta,\leavevmode\nobreak\ {\mathcal{L}}(X_{i},\eta)=1\big{\}}\Big{)}.$ (24)
We prove the validity of $\widehat{C}_{\alpha,n+j}^{\textnormal{ps}}$ in (23) below, whose proof is deferred to Appendix B.9.
Proposition 8.
For any selection rule ${\mathcal{L}}$, any $j\in[m]$ and any selection taxonomy $\mathfrak{S}$ such that $j\in\widehat{{\mathcal{S}}}_{\textnormal{ps}}$ and $\widehat{{\mathcal{S}}}_{\textnormal{ps}}\in\mathfrak{S}$, the prediction set $\widehat{C}_{\alpha,n+j}^{\textnormal{ps}}$ is a superset of the JOMI prediction set $\widehat{C}_{\alpha,n+j}$ defined in (8).
The procedure is summarized in Algorithm 4. For each $j$, the computation cost of $\widehat{C}_{\alpha,n+j}^{\textnormal{ps}}$ is $O(m)$.
Input: Calibration data ${\mathcal{D}}_{\textnormal{calib}}$; test data ${\mathcal{D}}_{{\textnormal{test}}}$; selection taxonomy $\mathfrak{S}$;
selective miscoverage level $\alpha$; preliminary micoverage level $\beta$;
selection rule ${\mathcal{S}}$; selection score function $S(\cdot,\cdot)$; nonconformity score function $V(\cdot,\cdot)$.
Compute $q_{1}$ and $q_{2}$ as in (4.3). Compute $\eta^{+}$ as the $(K+1)$-th order statistic of $\{S_{i}\}_{i=1}^{n}$, where $K=\lceil(1-\beta)(n+1)\rceil$. Compute $\eta^{+}$ as the $(K-1)$-th order statistic of $\{S_{i}\}_{i=1}^{n}$ for $j\in\widehat{\mathcal{S}}={\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})$ do
       Compute $\widehat{C}_{\alpha,n+j}^{\textnormal{ps}}$ as in (23).
end for
Output: $\{\widehat{C}_{\alpha,n+j}\}_{j\in\widehat{{\mathcal{S}}}}$.
Algorithm 4 JOMI for selection based on preliminary prediction sets
5 Application to drug discovery
In drug discovery, powerful prediction machines are increasingly used to guide the search of promising drug candidates. For such high-stakes decision-making, it is important to quantify the uncertainty in complex prediction machines (Svensson et al.,, 2017; Jin and Candès, 2023b, ; Laghuvarapu et al.,, 2024). Meanwhile, scientists will naturally focus on a subset of seemingly promising drugs.
Application scenarios.
We study two problems in drug discovery:
  1. 1.
    Drug property prediction (DPP), a classification problem where the binary outcome indicates whether a drug candidate binds to a pre-specified disease target, and the covariates are the (encoded) chemical structure of the drug compound.
  2. 2.
    Drug-target-interaction prediction (DTI), a regression problem where each sample is a pair of drug and disease target. The outcome of interest is a real-valued variable indicating the binding affinity of that pair. The covariates are the (concatenated) encoded structure of the drug and the disease.
The datasets and prediction models are based on the DeepPurpose library (Huang et al.,, 2020), with details in subsequent subsections. In each problem, we consider several realistic selection rules.
Selection rules.
We consider three types of realistic selection rules ${\mathcal{S}}$:
  1. (1)
    Covariate-dependent top-K selection: selecting drugs with highest predicted binding affinities $\widehat{\mu}(X_{n+j})$.
    1. (i)
      Top-K selection for test data. When the scientist has a fixed budget of investigating $K$ drug candidates in the next phase, one may select $K$ test samples with the largest values of $\widehat{\mu}(X_{j})$.
    2. (ii)
      Top-K selection with mixed data. When the scientist is to investigate other properties for promisingly active drug candidates in the next phase, they may select $K$ units in ${\mathcal{D}}_{\textnormal{calib}}\cup{\mathcal{D}}_{\textnormal{test}}$ with the highest predicted affinities.
    3. (iii)
      Calibration-referenced selection. The scientist may use the calibration data as reference and select test samples whose predicted activities are greater than the $K$-th highest in $\{\widehat{\mu}(X_{i})\}_{i\in{\mathcal{I}}_{\textnormal{calib}}}$.
  2. (2)
    Conformalized selection. The scientist might also obtain a subset of active drugs while controlling the FDR below some $q\in(0,1)$. In this case, $\widehat{\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})$ is the set of test drugs picked by the conformalized selection method in Jin and Candès, 2023b at FDR level $q\in(0,1)$.
  3. (3)
    Selection with constraints. The cost of developing a drug is a function $C(x)$, and one wants to select as many drugs with highest predictions $\widehat{\mu}(X_{n+j})$ as possible while ensuring the total cost is below some constant $C$, or maximize the total reward while controlling the total budget.
For brevity, each selection rule is reported for one application, while we defer others to Appendix A.
Evaluation metrics.
We evaluate the selection-conditional (mis)coverage via the consistent estimator
$\widehat{\textrm{Miscov}}=\frac{\sum_{j=n+1}^{n+m}\widehat{P}(j\in\widehat{\mathcal{S}},Y_{j}\notin\widehat{\mathcal{C}}_{j})}{\sum_{j=n+1}^{n+m}\widehat{P}(j\in\widehat{\mathcal{S}})},$
where $\widehat{P}$ is the empirical probability over all repeats. We also evaluate the size of prediction sets by averaging the cardinality of $\widehat{\mathcal{C}}_{j}$ for classification or length of $\widehat{\mathcal{C}}_{j}$ for regression over selected test units in all repeats. We also evaluate (but do not show in figures for brevity) the false coverage rate via $\widehat{\textrm{FCR}}=\widehat{E}\Big{[}\frac{\sum_{j=n+1}^{n+m}\mathds{1}\{j\in\widehat{\mathcal{S}},Y_{j}\notin\widehat{\mathcal{C}}_{j}\}}{1\vee|\widehat{\mathcal{S}}|}\Big{]},$ where $\widehat{E}$ is the empirical mean over all repeats.
5.1 Drug property prediction
We first consider the task of predicting the binary binding property of drug candidates. We use the HIV screening data in DeepPurpose with a total sample size of $n_{\textnormal{tot}}=41127$. The numerical features $X\in{\mathcal{X}}$ are encoded by Extended-Connectivity FingerPrints (Rogers and Hahn,, 2010, ECFP) which characterize topological properties of the drug candidates. A small-scale neural network $\widehat{\mu}\colon{\mathcal{X}}\to[0,1]$ is trained on randomly sampled $20\%$ of the entire dataset. Then, the remaining is randomly split into two equally-sized folds ${\mathcal{D}}_{\textnormal{calib}}$ and ${\mathcal{D}}_{\textnormal{test}}$. The exchangeability among data points is thus satisfied. Sampling of the training data and model training is independently repeated $10$ times; within each, we randomly split calibration/test data for $100$ times. This leads to $N=1000$ independent runs.
5.1.1 Top-K selection
We first consider the selection rule (i) top-K with test data. For a trained predictor $\widehat{\mu}$, we experiment with selecting (a) the highest $K$ values of $\widehat{\mu}(X_{j})$ and (b) the lowest $K$ values of $\widehat{\mu}(X_{j})$ among all test data. We observe similar patterns in (ii) top-K selection with both test and calibration data, as well as (iii) calibration-referenced selection. For brevity, we defer results for these selection rules to Appendix A.1.
For a confidence level $\alpha\in(0,1)$, we apply the vanilla conformal prediction and JOMI (both deterministic and randomized) to the selected units at level $\alpha$, with $\mathfrak{S}=2^{{\mathcal{I}}_{{\textnormal{test}}}}$. Since the selection size is fixed, valid selection-conditional coverage implies FCR control at level $\alpha$. We vary $\alpha\in\{0.1,0.2,\dots,0.9\}$ and $K\in\{20,100,1000,2000,5000,10000,15000\}$. We set $V$ as the APS score (Romano et al.,, 2020)
Figure 2 shows the empirical selection-conditional coverage $\widehat{\textnormal{Miscov}}$ on the left panel and the average prediction set size on the right panel, both at marginal miscoverage levels $\alpha\in\{0.1,0.8\}$. The solid (dashed) lines show the results when we select units with the highest (lowest) predicted affinities.
Figure 2: Empirical selection-conditional miscoverage (left) and prediction set size (right) in drug property prediction for vanilla conformal prediction (Vanilla_CP), JOMI and randomized JOMI (JOMI_rand) for test units whose $\widehat{\mu}(X_{n+j})$ are top-$K$ (highest, solid line) or bottom-$K$ (lowest, dashed line) among test units. The reds dashed line is the nominal miscoverage level $\alpha\in\{0.1,0.8\}$. The results are averaged over $N=1000$ runs.
The orange curves (Vanilla_CP) show that vanilla conformal prediction with APS score is over-confident for units with the lowest predicted affinities while under-confident for units with highest predicted affinities. In contrast, the purple (JOMI) and blue (JOMI_rand) curves both show valid coverage for our proposed methods. Also, using the APS score introduces visible gap between the actual coverage and $1-\alpha$ for JOMI due to distretization, which is made exact by its randomized version.
Interestingly, vanilla conformal prediction with the APS score yields zero-cardinality prediction sets for units with lowest predicted affinities with $\alpha=0.8$ (the orange dashed line in the right panel of Figure 2). This is because vanilla CP covers other test units with very high rate, and thus marginal coverage is guaranteed even with empty prediction sets for selected units. Of course, this is worrying if one cares more about these units at the bottom.
The selective behavior of these methods also depends on the choice of the nonconformity score $V$. Appendix A.2 shows the results for selection rules (i)-(iii) with the binary score. While our proposed methods consistently achieve valid coverage for selected units, with the binary score, vanilla CP is over-confident for units with highest predicted affinities yet under-confident for units with lowest predicted affinities. Also, the coverage gap of JOMI due to discretization is less visible with the binary score.
5.1.2 Conformalized selection
We then consider conformalized selection, where the focal test units are those believed to obey $Y=1$ with false discovery rate control at level $q\in(0,1)$. This problem was investigated in Bao et al., (2024) with no exact finite-sample coverage guarantees in theory (though their heuristic method performs reasonably in their empirical studies). We apply conformalized selection (Jin and Candès, 2023b, ) at FDR level $q\in\{0.2,\dots,0.9\}$ and $c_{i}\equiv 0.5$ to determine $\widehat{\mathcal{S}}$, and construct prediction intervals for units in $\widehat{\mathcal{S}}$ using either the APS score or the binary score. Experiments are repeated for $N=1000$ independent runs.
Figure 3 depicts the miscoverage $\widehat{\textnormal{Miscov}}$ and prediction interval lengths with nominal miscoverage level $\alpha\in\{0.1,0.8\}$ under various FDR levels $q$ and two choices of nonconformity score $V$.
Figure 3: Empirical selection-conditional miscoverage (left) and prediction set size (right) in drug property prediction for test units selected by conformalized selection (for finding $Y=1$ at target FDR levels $q\in\{0.1,0.2,\dots,0.9\}$) when applying vanilla conformal prediction (Vanilla_CP), JOMI (JOMI), randomized JOMI (JOMI_rand), the method of Bao et al., (2024) (Heuristic and Heuristic_rand). Results with the APS score are in solid lines, while those with the binary score are in dashed lines. Details are otherwise as Figure 2.
Vanilla CP is still less calibrated for selected units: it is over-confident with both scores for $\alpha=0.1$, while being over-confident with binary score and under-confident with APS score at $\alpha=0.8$. In contrast, JOMI and JOMI_rand achieves valid selection-conditional coverage in all scenarios. There is some gap for JOMI with the APS score due to discretization, but not for JOMI_rand or the binary score. We observe an even lower empirical FCR than the conditional miscoverage for our methods (so they achieve valid FCR control); this is because the selection set can sometimes be empty for small values of FDR level $q$. Finally, compared with the heuristic methods of Bao et al., (2024), our method usually achieves smaller prediction set sizes whereas their method seems overly conservative.
5.1.3 Selection with constraints
We now consider selecting units with highest predicted binding affinities within a total budget of subsequent development. In this case, ${\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})=\{j\in{\mathcal{D}}_{\textnormal{test}}\colon\widehat{\mu}(X_{j})\geq\widehat{\mu}_{[k]}\}$, where $\widehat{\mu}_{[k]}$ is the $k$-th order statistic of $\{\widehat{\mu}(X_{j})\}_{j\in{\mathcal{D}}_{\textnormal{test}}}$, with $k=\max\{k\colon\sum_{\ell=k}^{m}L_{[k]}\leq C\}$, and $\{L_{j}\}_{j\in{\mathcal{D}}_{\textnormal{test}}}$ are the costs.
[…]

### v3

@@ B.2.1 Selection based on calibration quantiles
Similar to the previous two cases, directly computing ^$\widehat{{\mathcal{R}}}_{\textnormal{cq}}$ reduces the computation complexity of the reference set to $O(\max(m,n))$. As a side note, Bao et al., (2024) also considered the selection rule based on calibration quantiles (as an instance of “calibration-assisted selection”), for which they provide asymptotically FCR control guarantees under additional assumptions.
Appendix C Additional experiment results
This section collects additional experimental results that are omitted in Section 5 of the main text. Supplementary Section C.1 shows results for DPP with selection rules (ii) and (iii) with the APS score. Supplementary Section C.2 shows results for DPP with selection rules (i)-(iii) with the binary score. Supplementary Section C.5 shows results for DTI with selection rules (ii) and (iii).
C.1 Additional results for APS score in drug property prediction
In this part, we present results for DPP when test units are selected via (ii) top-K in both calibration and test data (Supplementary Section C.1.1), and (iii) only calibration data as reference (Supplementary Section C.1.2).
C.1.1 Top-K selection in mixed sample
We use the same values of $(\alpha,K)$ and the same scheme of repeated experiments as in Section 5.1 of the main text, with $\mathfrak{S}=2^{[m]}$. Figure 9 shows that vanilla conformal prediction incurs a large coverage gap for selected units, while JOMI and its randomized version both achieve valid selection-conditional coverage. The coverage gap due to discretization is eliminated again by randomization. Randomization also reduces the prediction set sizes on the right panel, but again, with $\alpha=0.8$, we see the concerning issue that it suffices for vanilla CP to construct empty prediction sets for selected units to achieve marginal coverage.
C.1.2 Calibration-referenced selection
Now we select test units whose predicted affinities must be greater (or smaller) than the $K$-th largest (or smallest) predictions in the calibration set. We use the choices of $(\alpha,K)$ and the scheme of repeated experiments as in Section 5.1 of the main text with $\mathfrak{S}=2^{[m]}$.
C.2 Additional results for binary score in drug property prediction
In Figures 1112, and 13, we show the empirical selection-conditional coverage and prediction set sizes for selection rules (i)-(iii) when the prediction sets are constructed with the binary score $V(x,y)=y(1-\widehat{\mu}(x))+(1-y)\widehat{\mu}(x)$. We observe that while vanilla CP is either over-confident or under-confident, our proposed methods achieve valid coverage across all configurations. The coverage gap of the non-randomized JOMI is smaller with this specific score function compared with the APS score.
Figure 13: Empirical selection-conditional miscoverage (left) and prediction set size (right) in drug property prediction with the binary score, for test units that are (iii) greater than the $K$-th highest (or smaller than the $K$-th lowest) predicted affinities in the calibration data. Details are as Figure 10.
C.3 Additional results for empirical FCR in drug property prediction
In this part, we present empirical FCR results in drug property prediction, which complement the three tasks in Section 5 in the main text.
Figure 14 plots the empirical FCR in DPP in three top-K selection tasks with two conformity scores (APS and binary) at nominal miscoverage level $\alpha=0.1$. Since the selection set size is fixed, by our theory, selection-conditional coverage implies FCR control, as evidence from the results.
Figure 15 plots the empirical FCR when the units are selected by the conformal selection algorithm in Jin and Candès, 2023b . Here, the empirical FCR of JOMI is sill very close to the nominal conditional coverage level, because the selection set size is stable.
Finally, Figure 16 plots the empirical FCR when the units are selected by running the constrained optimization program. While the empirical FCR of JOMI is close to the nominal conditional coverage level, the FCR of vanilla CP with both conformity scores is lower than the selection-conditional coverage. This is because the selection set can sometimes be empty, and FCR is considered on this event. This also shows the limitation of FCR as an overly “optimistic” error metric: even if the FCR is low, on the “interesting” event that the selection set is nonzero, the coverage might still be very low.
Figure 14: Empirical false coverage rate across $N=1000$ independent runs of Vanilla_CP, JOMI, and JOMI_rand in top-K selection for drug property prediction. The $x$-axis is the nominal miscoverage level $\alpha=0.1$.
Figure 15: Empirical false coverage rate across $N=1000$ independent runs of Vanilla_CP, JOMI, and JOMI_rand in drug property prediction when units are selected by FDR-controlled conformal selection. The $x$-axis is the nominal FDR level in the selection algorithm.
Figure 16: Empirical false coverage rate across $N=1000$ independent runs of Vanilla_CP, JOMI, and JOMI_rand in drug property prediction when units are selected by running the optimization program. The $x$-axis is the nominal miscoverage level $\alpha\in\{0.1,\dots,0.9\}$.
C.4 Application to drug-target-interaction prediction
[…]